OpenStudy (anonymous):

What is the missing reason in the proof? Prove –(–y – x) – x = y Statements Reasons –(–y – x) – x = –[–y +(–x)] – x Definition of subtraction –[–y +(–x)] – x = y + x – x Opposite of a sum property y + x – x = y + x + (–x) ________________________ y + x + (–x) = y + [x + (–x)] Associative Property of Addition y + [x + (–x)] = y + 0 Additive Inverse Property y + 0 = y Additive Identity Property A. Definition of Subtraction B. Additive Inverse Property C. Additive Identity Property D. Opposite of a Sum Property

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

What is the missing reason in the proof? Prove –(–y – x) – x = y Statements Reasons –(–y – x) – x = –[–y +(–x)] – x Definition of subtraction –[–y +(–x)] – x = y + x – x Opposite of a sum property y + x – x = y + x + (–x) ________________________ y + x + (–x) = y + [x + (–x)] Associative Property of Addition y + [x + (–x)] = y + 0 Additive Inverse Property y + 0 = y Additive Identity Property

OpenStudy (anonymous):

A. Definition of Subtraction B. Additive Inverse Property C. Additive Identity Property D. Opposite of a Sum Property

OpenStudy (anonymous):

D

OpenStudy (anonymous):

ty